Find equivalent expressions to the following: xy/z, x/(yz), x/yz, (x/y)z, x/y/z

1
Expert's answer
2021-09-13T06:43:20-0400

"\frac{xy}{z},\frac{x}{yz},\frac{xz}{y},\frac{xz}{y},
\frac{x}{yz}"